OpenStudy (anonymous):

y=3^(2(x-3))+1 sketch and properly label the toolkit function, then perform the transformations in the correct order, and sketch and label the transformation function

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

this seems to be an application of some text specific method ....

OpenStudy (anonymous):

graphing the toolkit function

OpenStudy (anonymous):

and then do horizontal/vertical transformations

OpenStudy (amistre64):

y= 3^x shift right by 3 stretch horizontally by a factor of 2 then move up by 1 seems to be the process

OpenStudy (anonymous):

so would it be that i graph y=3^x

OpenStudy (amistre64):

start with y=3^x yes

OpenStudy (anonymous):

so (-2, 1/9), (-1, 1/3), (0, 1), (1, 3), 2,9

OpenStudy (amistre64):

those are points on 3^x yes

OpenStudy (amistre64):

now add 3 to your x values

OpenStudy (amistre64):

 at x=3, we should represent the value at x=0, since x-3 = 0 when x=3